### v1

# GUE correlators and large genus asymptotics

Abstract.
In this paper, we use a formula obtained in [5] to study certain asymptotic behaviors of GUE (Gaussian unitary ensemble) correlators. More precisely, we obtain large genus asymptotics of enumerations of ordinary graphs and ribbon graphs with 1 face.
1. Introduction
Large genus asymptotics of psi-class intersection numbers on the moduli space of stable algebraic curves have been studied in the literature. Liu and Xu found and proved [9] certain large genus asymptotics and rationality properties. Delecroix, Goujard, Zograf and Zorich (DGZZ) conjectured [4] the large genus uniform leading asymptotics of psi-class intersection numbers, which was improved and proved by Aggarwal in [1]. Guo and Yang [7] gave a new proof of the above-mentioned Liu-Xu’s results and DGZZ’s conjecture through estimates of the matrix-resolvent formula given in [2] regarding the KdV hierarchy.
In this paper, using the matrix-resolvent formula for the GUE (Gaussian unitary ensemble) correlators in [5] (which are closely related to enumerations of ribbon graphs), we will study certain large genus asymptotic behaviors and certain rationality properties. Our method is similar to [7]. In particular, we will obtain large genus asymptotics of enumerations of ordinary graphs and ribbon graphs with 1 face.
The GUE correlators of observables $\mathrm{tr}M^{i},i=1,2,\cdots$, are defined by
(1) $\langle\text{tr}M^{i_{1}}\cdots\text{tr}M^{i_{n}}\rangle_{c}(N):=\frac{\int_{\mathcal{H}(N)}\text{tr}M^{i_{1}}\cdots\text{tr}M^{i_{n}}e^{-\frac{1}{2}\text{tr}M^{2}}\mathop{}\!\mathrm{d}M}{\int_{\mathcal{H}(N)}e^{-\frac{1}{2}\text{tr}M^{2}}\mathop{}\!\mathrm{d}M},$
where $\mathcal{H}(N)$ denotes the space of $N\times N$ Hermitian matrices. It is known from [5] that the GUE partition function
(2) $Z_{N}(\textbf{s};\epsilon)=\frac{(2\pi)^{-N}\epsilon^{-1/12}}{\text{Vol}(N)}\int_{\mathcal{H}(N)}e^{-\text{tr}V(M)}\mathop{}\!\mathrm{d}M,$
where $\mathcal{H}(N)$ is the space of $N\times N$ Hermitian matrices and
(3) $V(M)=\frac{1}{2}M^{2}-\sum_{j\geq 1}s_{j}M^{j},\quad\textbf{s}=(s_{1},s_{2},\cdots),$
is a tau-function of the Toda lattice hierarchy, with the time variables defined by $t_{j}=s_{j+1}/\epsilon,j=0,1,2,\cdots$.
For $n\geq 1$, denote the generating series for the $n$-point correlators by
(4) $C_{n}(N;\lambda_{1},\cdots,\lambda_{n}):=\sum_{i_{1},\cdots,i_{n}=1}^{\infty}\frac{\langle\text{tr}M^{i_{1}}\cdots\text{tr}M^{i_{n}}\rangle_{c}(N)}{\lambda_{1}^{i_{1}+1}\cdots\lambda_{n}^{i_{n}+1}}.$
Using the matrix resolvent method, Dubrovin and Yang [5] obtained an explicit formula for $C_{n}(N;\lambda_{1},\cdots,\lambda_{n})$, which reads as follows
(5) $C_{1}(N;\lambda)=N\sum_{j\geq 1}\frac{(2j-1)!!}{\lambda^{2j+1}}({{}_{2}F}_{1}(-j,-N;2;2)-j\cdot{{}_{2}F}_{1}(1-j,1-N;3;2)),$
(6) $C_{n}(N;\lambda_{1},\cdots,\lambda_{n})=-\frac{1}{n}\sum_{\sigma\in S_{n}}\frac{R_{N}(\lambda_{\sigma(1)})\cdots R_{N}(\lambda_{\sigma(n)})-\delta_{n,2}}{\prod_{i=1}^{n}(\lambda_{\sigma(i)}-\lambda_{\sigma(i+1)})},\quad(n\geq 2),$
where $S_{n}$ denotes the symmetric group, $\sigma(n+1):=\sigma(1)$, ${{}_{2}F}_{1}(a,b;c;z)$ is the Gauss hypergeometric function
(7) ${{}_{2}}F_{1}(a,b;c;z)=\sum_{j=0}^{\infty}\frac{(a)_{j}(b)_{j}}{(c)_{j}}\frac{z^{j}}{j!}=1+\frac{ab}{c}\frac{z}{1!}+\frac{a(a+1)b(b+1)}{c(c+1)}\frac{z^{2}}{2!}+\cdots,$
and $R_{N}(\lambda)$, $N\geq 1$ denote the following matrix-valued series
(8) $R_{N}(\lambda):=\begin{pmatrix}1&0\\ 0&0\end{pmatrix}+N\sum_{j=0}^{\infty}\frac{(2j-1)!!}{\lambda^{2j+2}}\begin{pmatrix}(2j+1)A_{N,j}&-\lambda B_{N+1,j}\\ \frac{\lambda}{N}B_{N,j}&-(2j+1)A_{N,j}\end{pmatrix}$,
where
(9) $A_{N,j}:={{}_{2}F}_{1}(-j,1-N;2;2),\quad B_{N,j}:={{}_{2}F}_{1}(-j,1-N;1;2).$
Denote by $\mathrm{OG}_{g}(i_{1},\cdots,i_{n})$ the weighted number of connected ordinary graphs of genus $g$ with $n$ vertices of valencies $i_{1},\cdots,i_{n}$. We know from Euler’s formula that
(10) $i_{1}+\cdots+i_{n}=2g+2n-2.$
Then we have
(11) $\mathrm{OG}_{g}(i_{1},\cdots,i_{n})=\frac{1}{i_{1}!\cdots i_{n}!n!}\langle\text{tr}M^{i_{1}}\cdots\text{tr}M^{i_{n}}\rangle_{c}(1).$
Introduce a normalization of the enumeration of ordinary graphs by
(12) $\mathcal{G}^{\mathrm{OG}}_{i_{1},\cdots,i_{n}}(g):=\frac{i_{1}!\cdots i_{n}!n!}{(2g+2n-3)!!}\mathrm{OG}_{g}(i_{1},\cdots,i_{n}).$
We will prove in Section 2 the following theorems for the asymptotic formulas and the rationality of the normalized numbers of ordinary graphs.
Theorem 1.
For fixed $n\geq 1$ and fixed integers $i_{1},\cdots,i_{n-1}\geq 1$, we have
(13) $\lim_{g\to+\infty}\mathcal{G}^{\mathrm{OG}}_{i_{1},\cdots,i_{n-1},2g+2n-2-|i|}(g)=1,$
where $|i|=i_{1}+\cdots+i_{n-1}$.
Theorem 2.
For fixed $n\geq 1$ and fixed integers $i_{1},\cdots,i_{n-1}\geq 1$, there exists a rational function $R^{\mathrm{OG}}(g;i_{1},\cdots,i_{n-1})$ of $g$ that may depend on $n,i_{1},\cdots,i_{n-1}$, such that
(14) $\mathcal{G}^{\mathrm{OG}}_{i_{1},\cdots,i_{n-1},2g+2n-2-|i|}(g)=R^{\mathrm{OG}}(g;i_{1},\cdots,i_{n-1}).$
Theorem 1 and Theorem 2 yield the following corollary.
Corollary 1.
For fixed $n\geq 1$ and fixed integers $i_{1},\cdots,i_{n-1}\geq 1$, we have the following convergent expansion:
(15) $\mathcal{G}^{\mathrm{OG}}_{i_{1},\cdots,i_{n-1},2g+2n-2-|i|}(g)\sim\sum_{k=0}^{\infty}\frac{\mathcal{G}^{\mathrm{OG}}_{k}}{g^{k}},\quad(g\to\infty),$
where $\mathcal{G}^{\mathrm{OG}}_{0}=1$ and $\mathcal{G}^{\mathrm{OG}}_{k}=\mathcal{G}^{\mathrm{OG}}_{k}(n,i_{1},\cdots,i_{n-1}),k\geq 1$ are constants.
Denote by $a_{g}(i_{1},\cdots,i_{n})$ the weighted number of connected ribbon graphs of genus $g$ with $n$ vertices of valencies $i_{1},\cdots,i_{n}$ (see [5]). Then using Euler’s formula, we know that the number of faces of such a graph is $2-2g-n+\frac{1}{2}\sum_{j=1}^{n}i_{j}$. Now we focus on graphs with 1 face. We denote the number of ribbon graphs of this kind by $\mathrm{RG}_{g}(i_{1},\cdots,i_{n})$. Then Euler’s formula implies that
(16) $i_{1}+\cdots+i_{n}=4g+2n-2.$
In fact, $\mathrm{RG}_{g}(i_{1},\cdots,i_{n})$ is the first order coefficient of $N$ in the polynomial $\frac{1}{n!}\langle\text{tr}M^{i_{1}}\cdots\text{tr}M^{i_{n}}\rangle_{c}(N)$ (see [3, 8], cf. [5]), i.e.,
(17) $\langle\text{tr}M^{i_{1}}\cdots\text{tr}M^{i_{n}}\rangle_{c}=n!\mathrm{RG}_{g}(i_{1},\cdots,i_{n})N+O(N^{2}).$
We introduce the normalized numbers of ribbon graphs as follows:
(18) $\mathcal{G}^{\mathrm{RG}}_{i_{1},\cdots,i_{n}}(g):=\frac{n!}{2\cdot(4g+2n-5)!!}\mathrm{RG}_{g}(i_{1},\cdots,i_{n}).$
We will prove in Section 3 the following two theorems.
Theorem 3.
For fixed $n\geq 1$ and fixed integers $i_{1},\cdots,i_{n-1}\geq 1$, we have
(19) $\lim_{g\to+\infty}\mathcal{G}^{\mathrm{RG}}_{i_{1},\cdots,i_{n-1},4g+2n-2-|i|}(g)=1,$
where $|i|=i_{1}+\cdots+i_{n-1}$.
Theorem 4.
For fixed $n\geq 1$ and fixed integers $i_{1},\cdots,i_{n-1}\geq 1$, there exists a rational function $R^{\mathrm{RG}}(g;i_{1},\cdots,i_{n-1})$ of $g$ that may depend on $n,i_{1},\cdots,i_{n-1}$, such that
(20) $\mathcal{G}^{\mathrm{RG}}_{i_{1},\cdots,i_{n-1},4g+2n-2-|i|}(g)=R^{\mathrm{RG}}(g;i_{1},\cdots,i_{n-1}).$
Theorem 3 and Theorem 4 yield the following corollary.
Corollary 2.
For fixed $n\geq 1$ and fixed integers $i_{1},\cdots,i_{n-1}\geq 1$, we have the following convergent expansion:
(21) $\mathcal{G}^{\mathrm{RG}}_{i_{1},\cdots,i_{n-1},4g+2n-2-|i|}(g)\sim\sum_{k=0}^{\infty}\frac{\mathcal{G}^{\mathrm{RG}}_{k}}{g^{k}}\quad(g\to\infty),$
where $\mathcal{G}^{\mathrm{RG}}_{0}=1$ and $\mathcal{G}^{\mathrm{RG}}_{k}=\mathcal{G}^{\mathrm{RG}}_{k}(n,i_{1},\cdots,i_{n-1}),k\geq 1$ are constants.
Acknowledgements The author would like to thank Di Yang for his advice, suggestions of the questions, and helpful discussions. The author is also grateful to Jindong Guo for helpful discussions. Part of the work was done while the author was an undergraduate student of USTC; he thanks USTC for excellent working conditions.
2. Asymptotic Formula on Enumeration of Ordinary Graphs
In this section, we prove Theorem 1 and Theorem 2.
As in [7], we consider the rational function
(22) $P(\sigma;\lambda_{1},\dots,\lambda_{n})=\frac{1}{\prod_{q=1}^{n}(\lambda_{\sigma(q)}-\lambda_{\sigma(q+1)})}.$
The formal Laurent series of the whole right-hand side of (6) is independent of the ordering. Thus, following [7], we fix the choice of the expanding region to be $|\lambda_{1}|>\dots>|\lambda_{n}|$. We have the following lemma.
Lemma 1.
For each $\sigma\in S_{n}$, the Laurent expansion of the rational function $P(\sigma;\lambda_{1},\dots,\lambda_{n})$ around $\infty$’s within the region $|\lambda_{1}|>\dots>|\lambda_{n}|$ is given by
(23) $P(\sigma;\lambda_{1},\dots,\lambda_{n})=(-1)^{m(\sigma)}\sum_{j_{1},\dots,j_{n}\geq 0}\prod_{q=1}^{n}\lambda_{\sigma(q)}^{J_{\sigma,q}(j_{q})-J_{\sigma,q-1}(j_{q-1})-1}.$
Here,
(24) $m(\sigma):={\rm card}\left\{q\in\{1,\dots,n\}\mid\sigma(q+1)<\sigma(q)\right\},$
and
(25) $J_{\sigma,q}(j):=\left\{\begin{array}[]{ll}-j-1,&\sigma(q)<\sigma(q+1),\\ j,&\sigma(q)>\sigma(q+1),\end{array}\right$.
where the indices are considered cyclically modulo $n$.
For $n\geq 2$, $\underline{d}=(d_{1},\dots,d_{n})\in\mathbb{Z}^{\geq 0}$, $\sigma\in S_{n}$, and $q=1,\dots,n$, we define $K_{\underline{d},\sigma,q}:\left(\mathbb{Z}^{\geq 0}\right)^{n}\rightarrow\mathbb{Z}$ by
(26) $K_{\underline{d},\sigma,q}(\underline{j}):=d_{\sigma(q)}+J_{\sigma,q}(j_{q})-J_{\sigma,q-1}(j_{q-1}),$
where $\underline{j}=(j_{1},\dots,j_{n})\in\left(\mathbb{Z}^{\geq 0}\right)^{n}$, and $J_{\sigma,q}$ are given in (25).
For $n\geq 2$, $1\leq r\leq n-1$ being integers, we call a permutation $\sigma\in S_{n}$ with $\sigma(n)=n$ an (r,nr)(r,n-r)-permutation, if
(27) $n=\sigma(n)>\sigma(1)>\dots>\sigma(n-r+1)=1,$
(28) $1=\sigma(n-r+1)<\sigma(n-r+2)<\cdots<\sigma(n)=n$
(namely, firstly decreasing then increasing, often called unimodal permutations). The set of all permutations of this type is denoted by $S_{n,n}^{(r,n-r)}$. We also denote $S_{n,n}^{\{2\}}=\cup_{r=1}^{n-1}S_{n,n}^{(r,n-r)}$. The following combinatorial facts will be used later.
(29) ${\rm card}\,S_{n,n}^{(r,n-r)}=\binom{n-2}{r-1}$,
(30) ${\rm card}\,S_{n,n}^{\{2\}}=2^{n-2}.$
For $k\geq 0$, define a series of matrices $L_{k}$ by
(31) $L_{k}:=\left\{\begin{array}[]{ll}\begin{pmatrix}1&0\\ 0&0\end{pmatrix},&k=0,\\ \begin{pmatrix}(k-1)!!&0\\ 0&-(k-1)!!\end{pmatrix},&k\text{\,even\,and\,}k\geq 2,\\ \begin{pmatrix}0&-k!!\\ (k-2)!!&0\end{pmatrix},&k\text{\,odd and\,}k\geq 1.\end{array}\right$.
Then, by definition,
(32) $R_{1}(\lambda)=\sum_{k\geq 0}L_{k}\lambda^{-k}.$
For $k_{1},\cdots,k_{n}\geq 0$, we also denote by
(33) $l_{k_{1},\cdots,k_{n}}=\text{tr}(L_{k_{1}}\cdots L_{k_{n}}).$
We make the convention that $l_{k_{1},\cdots,k_{n}}=0$ if some of $k_{1},\cdots,k_{n}$ are less than or equal to 1-1. Note that $l_{k_{1},\cdots,k_{n}}$ is invariant under cyclic permutations of its indices.
With the above notations, we obtain an explicit formula on the normalized number of ordinary graphs by using Lemma 1 and expanding the right-hand side of (6), which is given by the following lemma.
Lemma 2.
For $n\geq 2$, $g\geq 0$ and $\underline{i}=(i_{1},\cdots,i_{n})\in(\mathbb{Z}^{\geq 1})^{n}$ satisfying $i_{1}+\cdots+i_{n}=2g+2n-2$, the following formula holds true:
(34) $\mathcal{G}^{\mathrm{OG}}_{\underline{i}}(g)=\frac{1}{(2g+2n-3)!!}\frac{1}{n}\sum_{\sigma\in S_{n}}(-1)^{m(\sigma)+1}\sum_{\underline{j}\in(\mathbb{Z}^{\geq 0})^{n}}l_{K_{\underline{i},\sigma,1}(\underline{j}),\cdots,K_{\underline{i},\sigma,n}(\underline{j})}.$
Due to our convention, for each $\sigma\in S_{n}$, the second sum on the right-hand side of (34) is a finite sum. Moreover, when $i_{1},\cdots,i_{n-1}$ are fixed and $g$ is sufficiently large, the number of terms in the sum is independent of $g$. Now we are ready to prove Theorem 1 and Theorem 2.
Proof of Theorem 1.
For the case $n=1$, we know from (5) that $\mathcal{G}^{\mathrm{OG}}_{2g}(g)=1$ for every $g\geq 1$. So we consider the case $n\geq 2$. For fixed integers $i_{1},\cdots,i_{n-1}\geq 1$, let $i_{n}=2g+2n-2-|i|$. From the definition (26) of $K_{\underline{i},\sigma,q}$, for each $\sigma\in S_{n}$ satisfying $\sigma(n)=n$, we have
(35) $K_{\underline{i},\sigma,n}\geq 2g+2n-2-|i|$
(36) $K_{\underline{i},\sigma,1}+\cdots+K_{\underline{i},\sigma,n}=2g+2n-2.$
For simplicity we write $\underline{K}=(K_{\underline{i},\sigma,1}(\underline{j}),\cdots,K_{\underline{i},\sigma,n}(\underline{j}))$. Then from the definition of $l_{i_{1},\cdots,i_{n}}$, (35) and (36), we have for $n\geq 1$:
(37) $\lim\limits_{g\to\infty}\frac{1}{(2g+2n-3)!!}l_{\underline{K}}=\left\{\begin{array}[]{ll}1,&\underline{K}=(0,0,\cdots,0,0,2g+2n-2),\\ -1,&\underline{K}=(1,0,\cdots,0,0,2g+2n-3),\\ 0,&\text{otherwise}.\end{array}\right$.
Now we need to respectively count the number of solutions to the two equations
(38) $\underline{K}=(0,0,\cdots,0,0,2g+2n-2),$
(39) $\underline{K}=(1,0,\cdots,0,0,2g+2n-3),$
for $\underline{j}\in(\mathbb{Z}^{\geq 0})^{n}$. It turns out that the number of solutions to (38) for $\underline{j}$ is
$\left\{\begin{array}[]{ll}i_{1},&\sigma\in S_{n,n}^{\{2\}},\\ 0,&\sigma\not\in S_{n,n}^{\{2\}},\end{array}\right$.
and the number of solutions to (39) for $\underline{j}$ is
$\left\{\begin{array}[]{ll}i_{1}-1,&\sigma\in S_{n,n}^{(1,n-1)},\\ i_{1},&\sigma\in S_{n,n}^{(r,n-r)}\text{ with }r\not=1,\\ 0,&\sigma\not\in S_{n,n}^{\{2\}}.\end{array}\right$.
Therefore, from (37), Lemma 2 and the above statement, we have
(40) $\lim\limits_{g\to\infty}\mathcal{G}^{\mathrm{OG}}_{i_{1},\cdots,i_{n-1},2g+2n-2-|i|}(g)$
$=$ $\lim\limits_{g\to\infty}\frac{1}{(2g+2n-3)!!}\frac{1}{n}\sum_{\sigma\in S_{n}}(-1)^{m(\sigma)+1}\sum_{\underline{j}\in(\mathbb{Z}^{\geq 0})^{n}}l_{K_{\underline{i},\sigma,1}(\underline{j}),\cdots,K_{\underline{i},\sigma,n}(\underline{j})}$
$=$ $\lim\limits_{g\to\infty}\frac{1}{(2g+2n-3)!!}\sum_{\sigma\in S_{n,n}^{\{2\}}}(-1)^{m(\sigma)+1}\sum_{\underline{j}\in(\mathbb{Z}^{\geq 0})^{n}}l_{K_{\underline{i},\sigma,1}(\underline{j}),\cdots,K_{\underline{i},\sigma,n}(\underline{j})}$
$=$ $\lim\limits_{g\to\infty}\frac{1}{(2g+2n-3)!!}\sum_{r=1}^{n-1}\sum_{\sigma\in S_{n,n}^{(r,n-r)}}(-1)^{r+1}\sum_{\underline{j}\in(\mathbb{Z}^{\geq 0})^{n}}l_{K_{\underline{i},\sigma,1}(\underline{j}),\cdots,K_{\underline{i},\sigma,n}(\underline{j})}$
$=$ $\lim\limits_{g\to\infty}\sum_{r=1}^{n-1}(-1)^{r+1}\binom{n-2}{r-1}(i_{1}-(i_{1}-\delta_{1,r}))$
$=$ $1.$
Here, we recall that since $i_{1},\cdots,i_{n-1}$ are fixed, the number of the terms in the right-hand side of (34) is a constant independent of $g$ for sufficiently large $g$. Then the theorem is proved. ∎
Now let us proceed to prove Theorem 2.
Proof of Theorem 2.
For $n=1$, the statement easily follows from (5). Now consider the case $n\geq 2$. For fixed integers $i_{1},\cdots,i_{n-1}$ and $i_{n}=2g+2n-2-|i|$, using (35), (36) and (33), one can see that for each $\sigma\in S_{n}$, $\frac{1}{(2g+2n-3)!!}l_{\underline{K}}$ is a rational function of $g$. Since the number of the terms in the right-hand side of (34) is independent of $g$ for sufficiently large $g$, we conclude from (34) the existence of a rational function $R^{\mathrm{OG}}(g;i_{1},\cdots,i_{n-1})$ such that $\mathcal{G}^{\mathrm{OG}}_{i_{1},\cdots,i_{n}}(g)=R^{\mathrm{OG}}(g;i_{1},\cdots,i_{n-1})$. ∎
3. Asymptotic Formula on Enumeration of Ribbon Graphs with 1 Face
In this section, we prove Theorem 3 and Theorem 4.
Define a series of matrix-valued polynomials $B_{k}$ for each $k\geq 0$ by
(41) $B_{k}=\left\{\begin{array}[]{ll}\begin{pmatrix}1&0\\ 0&0\end{pmatrix},&k=0,\\ \begin{pmatrix}N(k-1)!!\cdot\frac{1+(-1)^{(k-2)/2}}{k}&0\\ 0&-N(k-1)!!\cdot\frac{1+(-1)^{(k-2)/2}}{k}\end{pmatrix},&k\text{\,even and\,}k\geq 2,\\ \begin{pmatrix}0&-N(k-2)!!\\ (k-2)!!\cdot(-1)^{(k-1)/2}&0\end{pmatrix},&k\,\text{odd and\,}k\geq 1.\end{array}\right$.
For $k_{1},\cdots,k_{n}\geq 0$, we denote
(42) $b_{k_{1},\cdots,k_{n}}=[N^{1}]\text{tr}(B_{k_{1}}\cdots B_{k_{n}}).$
We make the convention that $b_{k_{1},\cdots,k_{n}}=0$ if some of $k_{1},\cdots,k_{n}$ are less than or equal to 1-1. Note that $b_{k_{1},\cdots,k_{n}}$ is invariant under cyclic permutations of its indices. With these definitions, we obtain an explicit formula for the normalized number of ribbon graphs by using Lemma 1 and expanding the right-hand side of (6), which is given by the following lemma.
Lemma 3.
For $n\geq 2$, $g\geq 0$ and $\underline{i}=(i_{1},\cdots,i_{n})\in(\mathbb{Z}^{\geq 1})^{n}$ satisfying $i_{1}+\cdots+i_{n}=4g+2n-2$, the following formula holds true:
(43) $\mathcal{G}^{\mathrm{RG}}_{\underline{i}}(g)=\frac{1}{2\cdot(4g+2n-5)!!}\frac{1}{n}\sum_{\sigma\in S_{n}}(-1)^{m(\sigma)+1}\sum_{\underline{j}\in(\mathbb{Z}^{\geq 0})^{n}}b_{K_{\underline{i},\sigma,1}(\underline{j}),\cdots,K_{\underline{i},\sigma,n}(\underline{j})}.$
By our convention, for each $\sigma\in S_{n}$ the second sum on the right-hand side of (43) is a finite sum. Moreover, when $i_{1},\cdots,i_{n-1}$ are fixed and $g$ is sufficiently large, the number of terms in the sum is independent of $g$. Now we are ready to prove Theorem 3 and Theorem 4.
Proof of Theorem 3.
For the case $n=1$, we know from (5) that $\mathcal{G}^{\mathrm{RG}}_{4g}(g)=1$ for every $g\geq 1$. Now we consider the case $n\geq 2$. For fixed integers $i_{1},\cdots,i_{n-1}\geq 1$, let $i_{n}=4g+2n-2-|i|$. From the definition (26) of $K_{\underline{i},\sigma,q}$, for each $\sigma\in S_{n}$ satisfying $\sigma(n)=n$, we have
(44) $K_{\underline{i},\sigma,n}\geq 4g+2n-2-|i|,$
(45) $K_{\underline{i},\sigma,1}+\cdots+K_{\underline{i},\sigma,n}=4g+2n-2.$
For simplicity we write $\underline{K}=(K_{\underline{i},\sigma,1}(\underline{j}),\cdots,K_{\underline{i},\sigma,n}(\underline{j}))$. Then from the definition of $b_{k_{1},\cdots,k_{n}}$, (44) and (45), we have for $n\geq 2$:
(46) $\lim\limits_{g\to\infty}\frac{1}{2\cdot(4g+2n-5)!!}b_{\underline{K}}=\left\{\begin{array}[]{ll}\frac{4g+2n-3}{2(4g+2n-2)}(1+(-1)^{n}),&\underline{K}=(0,0,\cdots,0,0,4g+2n-2),\\ -\frac{(-1)^{n}}{2},&\underline{K}=(0,0,\cdots,0,1,4g+2n-3),\\ -\frac{1}{2},&\underline{K}=(1,0,\cdots,0,0,4g+2n-3),\\ 0,&\text{otherwise}.\end{array}\right$.
Now we need to respectively count the number of solutions to the three equations
(47) $\underline{K}=(0,0,\cdots,0,0,4g+2n-2),$
(48) $\underline{K}=(0,0,\cdots,0,1,4g+2n-3),$
(49) $\underline{K}=(1,0,\cdots,0,0,4g+2n-3),$
for $\underline{j}\in(\mathbb{Z}^{\geq 0})^{n}$. It turns out that the number of solutions to (47) for $\underline{j}$ is
$\left\{\begin{array}[]{ll}i_{1},&\sigma\in S_{n,n}^{\{2\}},\\ 0,&\sigma\not\in S_{n,n}^{\{2\}},\end{array}\right$.
the number of solutions to (48) for $\underline{j}$ is
$\left\{\begin{array}[]{ll}i_{1}-1,&\sigma\in S_{n,n}^{(n-1,1)},\\ i_{1},&\sigma\in S_{n,n}^{(r,n-r)}\text{ with }r\not=n-1,\\ 0,&\sigma\not\in S_{n,n}^{\{2\}},\end{array}\right$.
and the number of solutions to (49) for $\underline{j}$ is
$\left\{\begin{array}[]{ll}i_{1}-1,&\sigma\in S_{n,n}^{(1,n-1)},\\ i_{1},&\sigma\in S_{n,n}^{(r,n-r)}\text{ with }r\not=1,\\ 0,&\sigma\not\in S_{n,n}^{\{2\}}.\end{array}\right$.
Therefore, from (46), Lemma 3 and the above statement, we have
(50) $\lim\limits_{g\to\infty}\mathcal{G}^{\mathrm{RG}}_{i_{1},\cdots,i_{n-1},4g+2n-2-|i|}(g)$
$=$ $\lim\limits_{g\to\infty}\frac{1}{2\cdot(4g+2n-5)!!}\frac{1}{n}\sum_{\sigma\in S_{n}}(-1)^{m(\sigma)+1}\sum_{\underline{j}\in(\mathbb{Z}^{\geq 0})^{n}}b_{K_{\underline{i},\sigma,1}(\underline{j}),\cdots,K_{\underline{i},\sigma,n}(\underline{j})}$
$=$ $\lim\limits_{g\to\infty}\frac{1}{2\cdot(4g+2n-5)!!}\sum_{\sigma\in S_{n,n}^{\{2\}}}(-1)^{m(\sigma)+1}\sum_{\underline{j}\in(\mathbb{Z}^{\geq 0})^{n}}b_{K_{\underline{i},\sigma,1}(\underline{j}),\cdots,K_{\underline{i},\sigma,n}(\underline{j})}$
$=$ $\lim\limits_{g\to\infty}\frac{1}{2\cdot(4g+2n-5)!!}\sum_{r=1}^{n-1}\sum_{\sigma\in S_{n,n}^{(r,n-r)}}(-1)^{r+1}\sum_{\underline{j}\in(\mathbb{Z}^{\geq 0})^{n}}b_{K_{\underline{i},\sigma,1}(\underline{j}),\cdots,K_{\underline{i},\sigma,n}(\underline{j})}$
$=$ $\lim\limits_{g\to\infty}\left(\Biggl(\left(\frac{4g+2n-3}{2(4g+2n-2)}(1+(-1)^{n})-\frac{(-1)^{n}}{2}\right)i_{1}-\frac{1}{2}(i_{1}-1)\Biggr)\right$.
$\qquad+\sum_{r=2}^{n-2}(-1)^{r+1}\binom{n-2}{r-1}\left(\frac{4g+2n-3}{2(4g+2n-2)}(1+(-1)^{n})-\frac{(-1)^{n}}{2}-\frac{1}{2}\right)i_{1}$
$\left.\qquad+(-1)^{n}\Biggl(\left(\frac{4g+2n-3}{2(4g+2n-2)}(1+(-1)^{n})-\frac{1}{2}\right)i_{1}-\frac{(-1)^{n}}{2}(i_{1}-1)\Biggr)\right)$
$=$ $1,$
which proves Theorem 3. ∎
Proof of Theorem 4.
For $n=1$, the statement easily follows from (5). Now consider the case $n\geq 2$. For fixed integers $i_{1},\cdots,i_{n-1}$ and $i_{n}=4g+2n-2-|i|$, using (44), (45) and (42), one can see that for each $\sigma\in S_{n}$, $\frac{1}{2\cdot(4g+2n-5)!!}b_{\underline{K}}$ is a rational function of $g$. Since the number of the terms in the right-hand side of (43) is independent of $g$ for sufficiently large $g$, we conclude from (43) the existence of a rational function $R^{\mathrm{RG}}(g;i_{1},\cdots,i_{n-1})$ such that $\mathcal{G}^{\mathrm{RG}}_{i_{1},\cdots,i_{n}}(g)=R^{\mathrm{RG}}(g;i_{1},\cdots,i_{n-1})$. ∎